# FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning

### v6

@@ 6 Conclusion
In future work, we plan to use the formular knowledge in FormulaReasoning to improve the numerical reasoning capabilities of LLMs. Possible approaches include enhancing reasoning abilities through knowledge-driven methods or reinforcement learning methods based on formula feedback.
@@ Appendix A Appendix
A.1 Appendix for Dataset Construction
A.1.1 Prompts for Formula Normalization
@@
A.1.2 Examples of Removed Questions
@@
A.1.3 Prompt for Semantics Based Merging
@@
A.1.4 Prompt for Dataset Translation
@@
A.2 Appendix for Experimental Setup
A.2.1 Prompt for Evaluated LLMs
@@ A.2.2 Formula Retriever
A.2.3 Data Augmentation for Supervised Fine-Tuned Models
@@
A.2.4 Preference Data Generation
@@
A.2.5 Implementation Details
We accessed DeepSeek-R1 through the DeepSeek API [^5] https://api-docs.deepseek.com/, Qwen series models through the Alibaba Bailian platform [^6] https://bailian.console.aliyun.com/, and other models through OpenRouter [^7] https://openrouter.ai/ with their default hyper-parameters. Qwen2.5-Math-7B/1.5B-Instruct were fine-tuned on NVIDIA GeForce RTX 3090 GPUs with Huggingface Transformers and Pytorch 2.3, where we used a learning rate of 1e-4 and a batch size of 8 to fine-tune them. For Qwen2.5-Math-7B-Instruct model, we fine-tune it with LoRA (Hu et al., 2021) with r=8, alpha=32. For Qwen2.5-Math-1.5B-Instruct model, we fine-tune it with full fine-tuning. We tested the models after training for 3 epochs which is enough for training. We set temperature=0 for direct answer generation and set the maximum output length to 1,024. In the case of CoT-SFT, which directly outputted formulas along with numerical values and units of parameters, we set temperature=0 to obtain the response with the highest probability and set the maximum output length to 512 to obtain formulas and parameters. Training Qwen2.5-Math-7B-Instruct(lora) and Qwen2.5-Math-1.5B-Instruct(full) in CoT-SFT used 3 and 8 hours, respectively.
A.3 Appendix for Experimental Results
A.3.1 English Version of FormulaReasoning
The evaluation results for LLMs with CoT prompts on the English FormulaReasoning dataset are presented in Table 5. Large-scale models (e.g., GPT-5 and DeepSeek-R1) achieved competitive results with accuracies exceeding 87% on both HoF and HeF metrics. However, substantial performance gaps existed for smaller-scale models, where architectures like Gemma-3n-E4B-it and Llama-3.3-8b-Instruct exhibited accuracies below 35%, while even mid-sized models like GLM-4.5-Air only reached around 73%. We found that Gemma-3n and Qwen3-235B showed performance drops compared to their results on the Chinese version due to instruction-following issues in the English version, where these models failed to output required units as specified in the prompt.
@@ A.3.1 English Version of FormulaReasoning
This significant variance suggests that while parameter scaling yields notable improvements, there remains considerable optimization potential for resource-efficient models. Future work should investigate architectural adaptations and training strategies specifically tailored for compact models, which are critical for practical deployment scenarios where computational resources are constrained. The observed performance spectrum underscores the importance of scale-aware evaluation in formula-based reasoning research.
@@
Model HoF test HeF test Avg. GLM-4.5-Air 72.64 74.42 73.53 ERNIE-4.5-21B-A3B 63.20 56.85 60.13 Gemma-3n-E4B-it 2.66 2.84 2.75 Llama-3.3-8b-Instruct 32.45 31.52 32.00 Mistral-Small-3.2-24b-Instruct 70.70 68.99 69.87 GPT-oss-20B 86.20 84.75 85.48 Gemini-2.5-Flash-Lite 75.30 76.49 75.88 Seed-OSS-36B-Instruct 75.30 71.58 73.50 GPT-4o 75.79 74.94 75.38 GLM-4-plus 75.79 68.73 72.37 Qwen3-235B-A22B-Thinking-2507 62.47 66.67 64.50 DeepSeek-R1 87.17 87.60 87.38 o3-mini 84.02 82.17 83.13 GPT-5 88.38 88.89 88.63
A.3.2 PoT Prompts
The results of GPT-4o using CoT compared to PoT are shown in Table 6. CoT consistently outperformed PoT. One possible reason was that the effectiveness of PoT could be influenced by a hybrid of the language and coding capabilities of LLMs, but that of CoT was only related to the language capabilities. Moreover, for LLMs with strong calculation capabilities, the advantage of using Python to perform numerical calculations might be eroded. Therefore, on FormulaReasoning, the bottleneck for current LLMs was mainly in their ability to apply formulas, rather than in calculation.
@@ A.3.2 PoT Prompts
A.3.3 PRM Scores
Tables 7 and 8 show the PRM scores in one-step and multi-step settings, respectively.
@@ A.3.3 PRM Scores
Model HoF test HeF test Avg. GLM-4.5-Air 0.9301 0.9436 0.9369 ERNIE-4.5-21B-A3B 0.7753 0.7582 0.7668 Gemma-3n-E4B-it 0.3695 0.3660 0.3678 Llama-3.3-8b-Instruct 0.2984 0.2408 0.2696 Mistral-Small-3.2-24b-Instruct 0.8576 0.8133 0.8355 GPT-oss-20B 0.5704 0.5203 0.5454 Gemini-2.5-Flash-Lite 0.5821 0.5818 0.5820 Seed-OSS-36B-Instruct 0.8766 0.8764 0.8765 GPT-4o 0.6150 0.5583 0.5867 GLM-4-plus 0.6728 0.6591 0.6660 Qwen3-235B-A22B-Thinking-2507 0.6502 0.6476 0.6489 DeepSeek-R1 0.7395 0.7393 0.7394 o3-mini 0.7537 0.7556 0.7547 GPT-5 0.8196 0.8249 0.8223
@@
Model HoF test HeF test Avg. GLM-4.5-Air 0.9336 0.9438 0.9387 ERNIE-4.5-21B-A3B 0.8695 0.8711 0.8703 Gemma-3n-E4B-it 0.7366 0.7360 0.7363 Llama-3.3-8b-Instruct 0.5507 0.5288 0.5398 Mistral-Small-3.2-24b-Instruct 0.9427 0.9438 0.9433 GPT-oss-20B 0.5738 0.5266 0.5502 Gemini-2.5-Flash-Lite 0.8621 0.8719 0.8670 Seed-OSS-36B-Instruct 0.8789 0.8779 0.8784 GPT-4o 0.8655 0.8444 0.8550 GLM-4-plus 0.8726 0.8738 0.8732 Qwen3-235B-A22B-Thinking-2507 0.7946 0.7996 0.7971 DeepSeek-R1 0.7603 0.7535 0.7569 o3-mini 0.8104 0.8032 0.8068 GPT-5 0.8196 0.8248 0.8222
Table 9: Accuracy of supervised fine-tuned models.
Model HoF test HeF test Avg.
Qwen2.5-Math-7B-Instruct 65.62 58.91 62.27
   + CoT-SFT 72.64 +7.02+7.02 65.63 +6.72+6.72 69.14 +6.87+6.87
   + DA 67.55 +1.93+1.93 62.53 +3.62+3.62 65.04 +2.77+2.77
Qwen2.5-Math-1.5B-Instruct 64.41 55.56 59.99
   + CoT-SFT 66.34 +1.93+1.93 53.75 1.81-1.81 60.05 +0.06+0.06
   + DA 66.59 +2.18+2.18 59.17 +3.61+3.61 62.88 +2.89+2.89
A.3.4 Supervised Fine-Tuned Models
Interestingly, by adding CoT-SFT, Qwen2.5-Math-7B-Instruct with a mere parameter count of 7B achieved a performance better than GPT-4o, although such a comparison might not be entirely fair. The results revealed that focusing on reasoning with CoT while delegating numerical calculations to a calculator greatly improved the performance of small models, given their limited calculation capacity. Data augmentation also lightly improved the reasoning capacity of small models.
@@ A.3.5 DPO
A.4 Error Analysis
FormulaReasoning posed challenges to existing models in terms of formula selection and numerical calculation, including unit calculation and arithmetic calculation. As illustrated in Figure 10(a), the incorrect use of formulas was the main type of error. Another type of error was parameter noise, as illustrated in Figure 10(b). In terms of the performance of the models, we found that the models that achieved higher accuracy were less affected by the parameter noise injected into the questions (Section 3.4). For example, among the questions to which DeepSeek-R1 did not provide a correct answer, approximately 14% were influenced by the parameter noise, while for Qwen-max, the number rose to 28%. This indicated that our question alteration imposed further requirements on the model’s ability to correctly apply formulas and perform accurate reasoning.

### v7

@@ 5 Conclusion
Future work will use FormulaReasoning’s formula knowledge to enhance LLM numerical reasoning via knowledge-driven or reinforcement learning methods. Moreover, given the domain-agnostic nature of our construction framework, we aim to extend our implementation to other scientific and engineering domains, such as chemistry and finance, to build a more comprehensive multi-disciplinary formula reasoning benchmark.
@@ Appendix A Appendix
A.1 Related Work
A.1.1 Numerical Reasoning Datasets
A.1.2 Numerical Reasoning Methods
Our experiments implemented and evaluated representative existing methods as baselines for FormulaReasoning, including zero-shot CoT prompting for LLMs ranging from 4B to more than 100B parameters. We trained a specialized formula retriever to be used in RAG. We divided the reasoning process into formula generation, parameter extraction, and numerical calculation, and used data augmentation to enhance fine-tuned small models. We derived preference data and performed DPO.
A.2 Appendix for Dataset Construction
A.2.1 Prompts for Formula Normalization
@@
A.2.2 Examples of Removed Questions
@@
A.2.3 Prompt for Semantics Based Merging
@@
A.2.4 Prompt for Dataset Translation
@@
A.2.5 Question Coupling
In this section, we provide a detailed walkthrough of the Question Coupling method using a concrete example. This method increased reasoning difficulty by coupling the original question with a generated auxiliary question through a mathematical expression. All the questions had been verified by Numbat to meet the constraints and be solvable.
Original Question
Consider the following original question from FormulaReasoning:
[…]
Answer: 32C
Step 1: Parameter Masking
We first identified mutable parameters in the question. In this example, we randomly selected the heat released (880 J) as the parameter to mask, denoted as $[X]$. We also designated the original answer (32C) as the second unknown parameter $[Y]$.
  • $[X]$: Heat released (880)
  • $[Y]$: Original temperature (32)
Step 2: Expression Construction
We constructed a linear expression involving $[X]$ and $[Y]$. To ensure numerical stability and reasonable magnitude, we adjusted the coefficients based on the order of magnitude of the parameters. In this example, we aligned the magnitudes and constructed the expression:
$E(X,Y)=50\cdot Y-1\cdot X$
Substituting the values:
$50\times 32-1\times 880=1600-880=720$
The result of the expression is 720.
Step 3: Auxiliary Question Generation
We then generated an independent physics question whose answer corresponds to the calculated result (720). We randomly assigned a unit (e.g., Newtons) to this value. Using GPT-4o, we generated the following auxiliary question:
Auxiliary Question: Ming is conducting a physics experiment. He needs to use a crane to lift a mass of 80 kg vertically upwards at a constant speed. Given that the local gravitational acceleration is 9.0 m/s2, what force does the crane need to exert to make the mass rise at a constant speed?
The answer to this auxiliary question is $F=mg=80\times 9.0=720$ N, which matches our target value.
Step 4: Final Compilation
Finally, we assembled a new hybrid question. We injected a distractor parameter (e.g., “calorific value of gas”) to further increase difficulty. The final question requires the model to solve the auxiliary question to find the value of the expression (i.e., 720 here), and then solve a system of equations in $X$ and $Y$ where one equation represents their relation in the original question and the other is, in this example, $720=E(X,Y)$.
[…]
A.3 Appendix for Experimental Setup
A.3.1 Evaluated LLMs
We conducted experiments across a diverse spectrum of models including cost friendly small-scale models including ERNIE-4.5-21B-A3B Baidu-ERNIE-Team (2025), Gemma-3n-E4B-it Gemma (2025), GPT-oss-20B (Agarwal et al., 2025), alongside high-cost LLMs such as GPT-4o (OpenAI, 2024), GLM-4-plus (GLM et al., 2024), Qwen3-235B-A22B (Qwen3-235B-A22B-Thinking-2507) (Yang et al., 2025), DeepSeek-R1 (DeepSeek-AI et al., 2025), o3-mini (OpenAI, 2025b) and GPT-5 (OpenAI, 2025a).
A.3.2 Prompt for Evaluated LLMs
@@ A.3.3 Formula Retriever
A.3.4 Data Augmentation for Supervised Fine-Tuned Models
@@
A.3.5 Preference Data Generation
@@
A.3.6 Implementation Details
We accessed DeepSeek-R1 through the DeepSeek API [^5] https://api-docs.deepseek.com/, Qwen series models through the Alibaba Bailian platform [^6] https://bailian.console.aliyun.com/, GLM-4-plus through ZHIPU–AI API [^7] https://open.bigmodel.cn/dev/api/normal-model/glm-4, and other models through OpenRouter [^8] https://openrouter.ai/ with their default hyper-parameters. We set temperature=0 for direct answer generation and set the maximum output length to 1,024. In the case of CoT-SFT, which directly outputted formulas along with numerical values and units of parameters, we set temperature=0 to obtain the response with the highest probability and set the maximum output length to 512 to obtain formulas and parameters. Training Qwen2.5-Math-7B-Instruct(lora) and Qwen2.5-Math-1.5B-Instruct(full) in CoT-SFT used 3 and 8 hours, respectively.
A.4 Appendix for Experimental Results
A.4.1 English Version of FormulaReasoning
The evaluation results for LLMs with CoT prompts on the English FormulaReasoning dataset are presented in Table 7. Large-scale models (e.g., GPT-5 and DeepSeek-R1) achieved competitive results with accuracies exceeding 87% on both HoF and HeF metrics. However, substantial performance gaps existed for smaller-scale models, where architectures like Gemma-3n-E4B-it exhibited poor accuracies. We found that Gemma-3n-E4B-it and Qwen3-235B-A22B showed performance drops compared to their results on the Chinese version due to instruction-following issues in the English version, where these models failed to output required units as specified in the prompt.
@@ A.4.1 English Version of FormulaReasoning
There was a general trend of performance decline on FormulaReasoning+. Notable drops were seen in models like GPT-4o (from 75.38% to 13.48%), o3-mini (from 83.13% to 68.38%), and GPT-5 (from 88.63% to 70.51%). This confirmed that the the extended FormulaReasoning+ successfully introduced greater complexity, providing a more rigorous benchmark for evaluating the reasoning capabilities of LLMs.
This substantial performance disparity indicates that although scaling parameters leads to marked improvements, there is still significant room for optimizing resource-efficient models. Future research should prioritize architectural and training optimizations for compact models, which are essential for resource-constrained environments. The performance variation observed across models highlights the critical need for scale-aware evaluation in formula-based reasoning.
@@
FormulaReasoning FormulaReasoning+ Model HoF HeF Avg. HoF HeF Avg. ERNIE-4.5-21B-A3B 63.20 56.85 60.13 7.39 9.26 8.33 Gemma-3n-E4B-it 2.66 2.84 2.75 2.71 2.65 2.68 GPT-oss-20B 86.20 84.75 85.48 53.45 57.67 55.56 GPT-4o 75.79 74.94 75.38 14.53 12.43 13.48 GLM-4-plus 75.79 68.73 72.37 17.73 20.11 18.92 Qwen3-235B-A22B 62.47 66.67 64.50 77.34 77.51 77.43 DeepSeek-R1 87.17 87.60 87.38 76.85 79.37 78.11 o3-mini 84.02 82.17 83.13 68.23 68.52 68.38 GPT-5 88.38 88.89 88.63 71.43 69.58 70.51
A.4.2 PoT Prompts
The results of GPT-4o using CoT compared to PoT are shown in Table 8. CoT consistently outperformed PoT. One possible reason was that the effectiveness of PoT could be influenced by a hybrid of the language and coding capabilities of LLMs, but that of CoT was only related to the language capabilities. Moreover, for LLMs with strong calculation capabilities, the advantage of using Python to perform numerical calculations might be eroded. The observed underperformance of PoT compared to CoT appears to also stem from the models’ faced challenge in accurate application of formulas within a programmatic context. Therefore, on FormulaReasoning, the bottleneck for current LLMs was mainly in their ability to apply formulas, rather than in calculation.
@@ A.4.2 PoT Prompts
A.4.3 PRM Scores
Tables 9 and 10 show the PRM scores in one-step and multi-step settings, respectively.
@@ A.4.3 PRM Scores
Model HoF test HeF test Avg. ERNIE-4.5-21B-A3B 0.7753 0.7582 0.7668 Gemma-3n-E4B-it 0.3695 0.3660 0.3678 GPT-oss-20B 0.5704 0.5203 0.5454 GPT-4o 0.6150 0.5583 0.5867 GLM-4-plus 0.6728 0.6591 0.6660 Qwen3-235B-A22B 0.6502 0.6476 0.6489 DeepSeek-R1 0.7395 0.7393 0.7394 o3-mini 0.7537 0.7556 0.7547 GPT-5 0.8196 0.8249 0.8223
@@
Model HoF test HeF test Avg. ERNIE-4.5-21B-A3B 0.8695 0.8711 0.8703 Gemma-3n-E4B-it 0.7366 0.7360 0.7363 GPT-oss-20B 0.5738 0.5266 0.5502 GPT-4o 0.8655 0.8444 0.8550 GLM-4-plus 0.8726 0.8738 0.8732 Qwen3-235B-A22B 0.7946 0.7996 0.7971 DeepSeek-R1 0.7603 0.7535 0.7569 o3-mini 0.8104 0.8032 0.8068 GPT-5 0.8196 0.8248 0.8222
@@ A.4.4 DPO
A.5 Error Analysis
FormulaReasoning posed challenges to existing models in terms of formula selection and numerical calculation, including unit calculation and arithmetic calculation. As illustrated in Figure 10(a), the incorrect use of formulas was the main type of error. Another type of error was parameter noise, as illustrated in Figure 10(b). In terms of the performance of the models, we found that the models that achieved higher accuracy were less affected by the parameter noise injected into the questions (Section 2.4). For example, among the questions to which DeepSeek-R1 did not provide a correct answer, approximately 14% were influenced by the parameter noise. This indicated that our question alteration imposed further requirements on the model’s ability to correctly apply formulas and perform accurate reasoning.